40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 6 03:30:17 2015, ericq, Update, LSC, DRFPMI Progress 
    Reply  Thu Oct 8 04:48:50 2015, ericq, Update, LSC, DRFPMI Progress 
       Reply  Thu Oct 8 13:13:20 2015, Koji, Update, LSC, DRFPMI Progress 
          Reply  Thu Oct 8 14:14:50 2015, ericq, Update, LSC, DRFPMI Progress 
             Reply  Thu Oct 8 16:48:23 2015, Koji, Update, LSC, DRFPMI Progress 
             Reply  Thu Oct 8 21:35:49 2015, rana, Update, LSC, DRFPMI Progress 
                Reply  Fri Oct 9 09:22:38 2015, ericq, Update, LSC, DRFPMI Progress oct8_allRF.pdf
                   Reply  Fri Oct 9 11:24:06 2015, Jenne, Update, LSC, DRFPMI Progress 
Message ID: 11669     Entry time: Tue Oct 6 03:30:17 2015     Reply to this: 11671
Author: ericq 
Type: Update 
Category: LSC 
Subject: DRFPMI Progress 

[ericq, Gautam]

Highlight of the night: the DRFPMI was held at arm powers > 110 for 20 seconds. ALS feedback was still running though, but so was some nonzero REFL11 AO path action.

In short, time was spent finding the right FM trigger settings to keep the DRMI locked while CARM is fluctuating through resonance, what CARM offset to acquire DRMI lock at, order of operations of turning on AO / turning up overall CARM gain, etc. 

Sadly, for the past hour or so, the DRMI has refused to stay locked for more than ~20 seconds, so I haven't been able to push things much further. This is a shame, since I'm very nearly at the equivalent point in the PRFPMI locking script where the ALS control is turned off completely. 

ELOG V3.1.3-